OpenStudy (anonymous):

Calc III homework probs (with a friend) Calc III homework probs (with a friend) @Mathematics

OpenStudy (anonymous):

f(x,y)=x^3(y)-xy+y^3(x) find all crit points and classify each as a relative max, min, or saddle point using the second partials test.

OpenStudy (anonymous):

find the first partials first right? So you have: \[f_x(x,y)=3yx^2-y+y^3; f_y(x,y)=x^3-x+3xy^2\] Set those to zero. \[0=y(3x^2-1+y^2); x(x^2-1+3y^2)=0\] So if x=0, then y=0 holds. Then we have:\[3x^2-1+y^2=0 \implies x^2=\frac{1}{3}(1-y^2) \implies x=\pm \sqrt{\frac{1}{3}(1-y^2)}\] From this we can see that if that holds then: \[\frac{1}{3}(1-y^2)-1+3y^2=0 \implies -\frac{2}{3}+\frac{5}{3}y^2=0 \implies 5y^2=2 \implies y=\pm \sqrt{\frac{2}{5}}\] Which means that: \[x=\pm \sqrt{\frac{1}{5}}\] Then calculate the derivatives and plug those in. Done.

OpenStudy (anonymous):

\[y=0; x=\pm 1 \] Also works.

OpenStudy (anonymous):

x=0, y=pm 1 works

OpenStudy (anonymous):

Find all critical points of the function and determine the absolute max and mins on the region R. \[\frac{24xy}{(x^2+4)(y^2+9)}\]

OpenStudy (anonymous):

R = {(x,y):0<=x<=1 and 0<=y<=5}

OpenStudy (anonymous):

Just find the critical points like normal and then make sure they fall in the range r its not hard.